Question: Find the function composition of: f o g(x) given f(x) = x^2 - 1 and g(x) = x^2 - 4
What I've done so far:
(f o g)(x) = f(g(x)) (setting up the function)
f(x^2 - 4) (inputted g into the function)
x^2(x^2 - 4) - 1 (inputted f)
x^4 - 4x^2 - 1 (simplified)
I'm not sure if this is right and I'm not sure if I can simplify further by factoring or if this is even the end result.

Answer by stanbon(57361)

Find the function composition of: f o g(x) given f(x) = x^2 - 1 and g(x) = x^2 - 4
What I've done so far:
(f o g)(x) = f(g(x)) (setting up the function)
f(x^2 - 4) (inputted g into the function)
----
Note: The "x" in f is now (x^2-4).
---
= (x^2-4)^2-1
------
= x^4-8x^2+16-1
---
= x^4-8x^2+15
===================
Note: You could factor that but that is not
what the question asks you to do.
===========================================
Cheers,
Stan H.